"How many pounds of cashews valued at $2.00 per pound must be mixed with 30 pounds of peanuts valued at 80cents per pond to produce a mixture of $1.25 per pound?"
Answer by checkley77(12844)

2X+.80*30=1.25(30+X)
2X+24=37.5+1.25X
2X-1.25X=37.5-24
.75X=13.5
X=13.5/.75
X=18 POUNDS OF CASHEWS IS USED.
PROOF:
2*18+.8*30=1.25(30+18).
36+24=1.25*48
60=60